OpenStudy (anonymous):

when the price of a certain commodity is p dollars per unit, costumers demand x hundred units of the commodity, where .5x^2 + 2px +3p^2= 249. How fast is the demand x changing with respect to time when the demand is 6 units, the price is $7 per unit and is decreasing at the rate of 10 cents per month?

OpenStudy (perl):

you didnt give me a medal for my last question :/

OpenStudy (anonymous):

ill do that right now

OpenStudy (anonymous):

can you help me with this

OpenStudy (perl):

ok lets try implicit deriv.

OpenStudy (perl):

.5x^2 + 2px +3p^2= 249. .5*2x*dx/dt + 2*dp/dt*x + 2*p*dx/dt + 3*2*p*dp/dt = 0

OpenStudy (anonymous):

after this whats next

OpenStudy (perl):

now we can plug in

OpenStudy (anonymous):

can you tell me what to plug in

OpenStudy (perl):

when the price of a certain commodity is p dollars per unit, costumers demand x hundred units of the commodity, How fast is the demand x changing with respect to time when the demand is 6 units, the price is $7 per unit and is decreasing at the rate of 10 cents per month? x = 6 hundred units = 600 p = 7 dp/dt = -.10 .5*2x*dx/dt + 2*dp/dt*x + 2*p*dx/dt + 3*2*p*dp/dt = 0

OpenStudy (perl):

does it say demand is 6 hundred units?

OpenStudy (anonymous):

no it say demand is 6 units

OpenStudy (perl):

ok

OpenStudy (perl):

lets just plug in x = 6

OpenStudy (anonymous):

did you get 27

OpenStudy (perl):

x = 6 hundred units p = 7 dp/dt = -.10 .5*2*(6)*dx/dt + 2*(-.10)*6 + 2*7*dx/dt + 3*2*7*(-.10) = 0

OpenStudy (perl):

yes I got dx/dt = .27, then multiply by 100, so 27

OpenStudy (anonymous):

ok

OpenStudy (perl):

are these answers correct ?

OpenStudy (perl):

did we get the last one correct ?

OpenStudy (anonymous):

can you help me with one more

OpenStudy (perl):

need to know if we are doing them correctly

OpenStudy (anonymous):

we are

OpenStudy (perl):

ok so then the correct answer is 27